how to solve difficult problems of complex numbers with trigonometry??? can u explain the basic funda to apply in complex numbers questions???? if possible, please explain in brief reasoning behind them....


thank you.

Grade:12

1 Answers

piyush garg
14 years ago

for example take a problem Z= cos(5 degree)+isin(5degree) you have to evalute Z^6 then we have to

use euler form Z=e^i(5) and Z^6=e^i(30)=cos(30)+isin(30)

which can be evaluated by use of trignometry.

i take very simple problem.you will take some other problems